OpenStudy (anonymous):

For what value of b will the line y= -2x + b be tangent to the parabola y= 3x^2 + 4x - 1?

OpenStudy (anonymous):

i hope this is a calculus course

OpenStudy (anonymous):

the derivative of \(y=3x^2+4x-1\) is \(y'=6x+4\)

OpenStudy (anonymous):

Not a calculus course :S

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

ok then \(y\) has to satisfy two equations \[y=-2x+b\] and \[y=3x^2+4x+1\] we can set \[3x^2+x+1=-2x+b\] and solve

OpenStudy (anonymous):

we get \[3x^2+3x+(1-b)=0\]

OpenStudy (anonymous):

in order for the line to be tangent, it must touch and not cross that means the above equation can have only one solution that means the discriminant must be zero

OpenStudy (anonymous):

Hm, the answers says -4

OpenStudy (anonymous):

all my algebra is wrong that is why lets start again

OpenStudy (anonymous):

\[3x^2+4x+1=-2x+b\] \[3x^2+6x+(1-b)=0\]

OpenStudy (anonymous):

now \[a=3,b=6,c=(1-b)\]

OpenStudy (anonymous):

\[b^2-4ac=36-12(1-b)=0\] \[36-12+12b=0\] \[24+12b=0\] \[24=-12b\] \[b=-2\]damn still wrong

OpenStudy (jdoe0001):

picking up from @satellite73 's line \(\bf y = 3x^2+4x-1\qquad tangent\quad \implies y = -2x+b\\ \quad \\ \quad \\ 3x^2+4x-1 \quad = \quad -2x+b\implies 3x^2+6x-1-b=0\\ \quad \\ 3x^2+6x-(1+b)=0\implies 3x^2+6x+(-1-b)=0\\ \quad \\ \textit{as satellite73 said, let us equate the discriminant to 0}\\ \quad \\ b^2-4ac = 0\implies (6)^2-4(3)(-1-b)=0\implies 36-12(-1-b)=0\\ \quad \\ 36+12+12b=0\implies 48+12b=0 \implies 12b=-48\implies b = -4\)